1) You receive $5,000 at the end of every year for three years. What is the present value of these receipts if you earn eight percent compounded annually

2)What amount will be accumulated in four years if $10,000 is invested today at six percent interest compounded annually?
Answer #1
1) Present value of the receipts
Year Cash Flow PV Factor PV Of Cash Flow
a b c=1/1.08^a d=c*b
1 $         5,000 0.925925926 $             4,629.63
2 $         5,000 0.85733882 $             4,286.69
3 $         5,000 0.793832241 $             3,969.16
Present value $           12,885.48
2)
Future value = PV * (1+r)^n
Where,
PV = Present value
t= interest rate
n= number of years
=$10000*(1+0.06)^4
=$12624.77
amount will be accumulated in four years =$12624.77
Kindly give a positive rating if you are satisfied with this solution and please ask if you have any queries.